OpenStudy (anonymous):

Medal and Fan! An inequality is shown below: -np-4 < or equal to 2(c-3) Which of the following solves for n? A) n < or equal to - 2c-2 over p B) n < or equal to - 2c-10 over p C) n > or equal to - 2c-2 over p D) n > or equal to - 2c-10 over p

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

I know its not C or D

OpenStudy (misty1212):

\[-np-4 \leq 2(c-3)\]\[-np-4\leq 2c-6\]\[-np\leq 2c-2\]

OpenStudy (anonymous):

So we would than divide by p?

OpenStudy (misty1212):

then divide by \(-p\) but you cannot solve it because you do not know if \(p\) is a positive or negative, but maybe it is supposed to be positive because it is called \(p\)

OpenStudy (anonymous):

So A :D

OpenStudy (misty1212):

if p is positive then dividing by \(-p\) means you have to change the inequality

OpenStudy (misty1212):

\[n\geq \frac{2-2c}{p}\]

OpenStudy (anonymous):

Yes but in each of my choices there is a new negative sign that is before every single fraction :3

OpenStudy (anonymous):

So if we divided by -P it would still make the entire fraction negative correct :3

OpenStudy (misty1212):

\[n\geq -\frac{2c-2}{p}\]

OpenStudy (anonymous):

Do we flip the sign each time?

OpenStudy (anonymous):

oh thats right when its a negative we do correct?

OpenStudy (misty1212):

if you think \(-p\) is negative then you flip the sign your math teacher is not very smart though, because it could be that \(-p\) is positive, like if \(p=-2\) you cannot solve an inequality with variable coefficients like that unless it says \(p>0\) your math teacher should be ashamed of his learning